OpenStudy (anonymous):

What is the third-degree polynomial function such that f(0) = -24 and whose zeros are 1, 2, and 3?

OpenStudy (anonymous):

From the given data, we know that\[f)x)=k(x-1)(x-2)(x-3)\]To find k, we consider the given information about the value of the function at x=0\[f(0)=24=k(-1)(-2)(-3)=-6k \implies k=-4 \]\[\implies f(x)=-4(x-1)(x-2)(x-3)\]

OpenStudy (anonymous):

Depending on your author or teacher, you might have to multiply the factors out to put the answer in the desired form.

OpenStudy (phi):

I would tweak Animal's answer, as he solved for f(0)= 24 and you want f(0)= -24. That means k = 4 (not -4) can you multiply out 4(x-1)(x-2)(x-3) ? As a first step, I would multiply (x-1)(x-2). Can you do that ?